OpenStudy (anonymous):

solve each equation. show work please.its a fraction just dnt know how to type on here c/-4 - 8 = -42

OpenStudy (anonymous):

c/(-4-8)=42 c/-12=42 -12/1*c/-12=42*-12 c=-504

jhonyy9 (jhonyy9):

c+32=168 c=136